A farmer plans to enclose a rectangular region, using part of his barn for one side and fencing for the other three sides. If the side parallel to the barn is to be twice the length of an adjacent side, and the area of the region is to be 128 ft squared, how many feet of fencing should be purchased?
Answer by nerdybill(7384)

A farmer plans to enclose a rectangular region, using part of his barn for one side and fencing for the other three sides. If the side parallel to the barn is to be twice the length of an adjacent side, and the area of the region is to be 128 ft squared, how many feet of fencing should be purchased?
.
Draw a diagram of the situation -- it is important so you can visualize what is going on.
.
Let x = measure of sides perpendicular to barn
then
2x = measure of side parallel to barn
.
area = x(2x) = 2x^2
128=2x^2
64 = x^2
8 ft = x (measure of side that is perpendicular to the barn)
2x = 16 ft (measure of side that is parallel to the barn)
.
fencing required is the perimeter.
perimeter = x+x+2x = 4x = 4(8) = 32 ft